Suppose that the risk-free rate is 3% and that the market risk premium is 8%.

What is the required return on the market? Round your answer to two decimal places. %

What is the required return on a stock with a beta of 1.2? Round your answer to two decimal places. %

What is the required return on a stock with a beta of 0.4? Round your answer to two decimal places.
Answer #1

Rf = Risk free rate =3%

Rm – Rf = Market Risk Premium = Return on market – Risk Free rate = 8%

So, Return on market = 8% + risk free rate

= 8 + 3

= 11%

As per capital asset pricing model, Required rate of return is given by

Re = Rf + ( Rm – Rf ) x Beta

= 3 + 8 x 1.20

= 3 + 9.6

= 12.6%

For situation 2

Re = 3 + 8 x 0.40

= 3 + 3.2

= 6.2%